Circular No. 8691

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2006 E1 (McNAUGHT) Further to IAUC 8688, R. H. McNaught reports that images taken

with the Uppsala Schmidt telescope by G. J. Garradd on Mar. 18.8 UT show this comet to be slightly diffuse with an extension ∼ 5′′ long to the northwest; additional images taken by McNaught on Mar. 23.7 show a similiar appearance. Additional astrometry, the following preliminary orbital elements, and an ephemeris appear on MPEC 2006-F28.

T = 2006 Dec. 14.954 TT  ω = 230.^{o}496
Ω = 94.949 }2000.0
q = 6.10616 AU  i = 82.985

SUPERNOVAE 2006ax, 2006ay, 2006az Three new apparent supernovae have been reported from unfiltered

CCD images: 2006ax and 2006ay by D. R. Madison, N. J. Ponticello, and W. Li (LOSS/KAIT; cf. IAUC 8680), and 2006az by J. Newton, L. Cox, and T. Puckett (0.35-m reflector at Portal, AZ; cf. IAUC 8664).

SN  2006 UT  α_{2000}  δ_{2000}  Mag.  Offset
2006ax Mar. 20.37 11^{h}24^{m}03.^{s}46 −12^{o}17′29′′.2 17.1 51′′ E, 18′′ N
2006ay Feb. 24.56 15 59 12.76 +20 44 27.9 19.0 12′′.2 E, 22′′.7 S
2006az Mar. 23.39 12 12 14.68 +56 10 45.8 17.1 1′′.7 E, 5′′.8 N

Additional approximate magnitudes for 2006ax in NGC 3663: Feb. 10.39 UT, [18.5 (KAIT); Mar. 21.15, 16.5 (S. Gonzalez, Las Campanas Observatory, 1.0-m Swope telescope). N. Tokimasa, H. Naito, and H. Yamaoka report that low-resolution spectroscopy obtained on Mar. 21.7 shows 2006ax to be a type-Ia supernova a few days before maximum light (additional details given on CBET 437). The offset for 2006ay above is with respect to the nucleus of the southeast component of UGC 10116 (a group of six galaxies also known as ‘Seyfert’s Sextet’). Additional approximate magnitudes for 2006ay: 2002 July 7.30, [19.8; 2006 Feb. 10.56, [18.8; Mar. 21.26, 20 (V -band image taken by Gonzalez). Nothing is visible at the location of 2006ay on the Digitized Sky Survey (limiting red mag ∼ 21; see also CBET 436). Additional magnitudes for 2006az in NGC 4172 from Puckett: Feb. 8, [19.0; Mar. 24.29, 16.3.

2006 March 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2006 CBAT  Daniel W. E. Green